# Crosstalk-Robust Quantum Control in Multimode Bosonic Systems

### v1

@@ Appendix B Ancilla gate decomposition for efficient experimental calibration
Figure 8: Infidelity of $\hat{X}_{\pi}$ and $\hat{X}_{\pi/2}$ as a function of detuning. The blue curves plot results obtained from the developed QOC pulses, while the red ones represent infidelities reported in Ref. 68. Note that the data below the gray line are not available in the reference.
Figure 9: Optimization results for the $\hat{X}_{\pi/2}$ gate. (a) Average infidelities in Eq. (19) versus pulse durations for both closed (dashed line) and open (solid line) systems. The optimal duration selected from the range considered for this optimization is 20 ns (indicated by a star). (b) $I$ and $Q$ quadratures of the pulse envelopes for the selected 20-ns pulse from (a). (c) The frequency components of the pulse envelopes, corresponding to the 20-ns pulse. (d) Comparative analysis of the 20-ns pulses, from both QOC (blue) and DRAG (red) pulses, in terms of their infidelities as a function of detuning. The average and standard deviation of the infidelities are tabulated correspondingly by considering a uniform distribution of $\rho(\delta)$ for simplicity. The result shows that the optimized pulse is more robust than the commonly used DRAG pulse. [Parameters: $\delta_{i}/2\pi\in\{\pm 1,\pm 3,\pm 4,\pm 7,\pm 10\}$ MHz, $K_{\text{q}}/2\pi=-200$ MHz, $T_{1}=50$ $\mu$s, $T_{\phi}=50$ $\mu$s]
Appendix C Supplement information on frequency-robust control of a transmon
C.1 Comparison of pulse robustness with recent literature
Here, we compare the robustness of the pulses derived from our optimization procedure with those reported in Ref. 68. For a fair comparison, we take the same parameters: transmon anharmonicity $K_{\text{q}}/2\pi=-260$ MHz, detuning range $\mathbb{D}/2\pi=[-4,2]$ MHz, durations for $\hat{X}_{\pi}$ and $\hat{X}_{\pi/2}$ gates as $70$ ns and $80$ ns, and no decoherence error. The comparison is illustrated in Fig. 8, where the blue and red curves represent our results and those from Ref. 68, respectively. Within the chosen range of detuning, our results consistently exhibit significantly lower infidelity.
C.2 Optimized 20-ns pulse for $\hat{X}_{\pi/2}$ gate

### v2

@@ Appendix D Ancilla gate decomposition for efficient experimental calibration
Appendix E Supplement information on frequency-robust control of a transmon
E.1 Optimized 20-ns pulse for $\hat{X}_{\pi/2}$ gate
Figure 8: Optimization results for the $\hat{X}_{\pi/2}$ gate. (a) Average infidelities in Eq. (12) versus pulse durations for both closed (dashed line) and open (solid line) systems. The optimal duration selected from the range considered for this optimization is 20 ns (indicated by a star). (b) $I$ and $Q$ quadratures of the pulse envelopes for the selected 20-ns pulse from (a). (c) The frequency components of the pulse envelopes, corresponding to the 20-ns pulse. (d) Comparative analysis of the 20-ns pulses, from both QOC (blue) and DRAG (red) pulses, in terms of their infidelities as a function of detuning. The average and standard deviation of the infidelities are tabulated correspondingly by considering a uniform distribution of $\rho(\delta)$ for simplicity. The result shows that the optimized pulse is more robust than the commonly used DRAG pulse. [Parameters: $\delta_{i}/2\pi\in\{\pm 1,\pm 3,\pm 4,\pm 7,\pm 10\}$ MHz, $K_{\text{q}}/2\pi=-200$ MHz, $T_{1}=50$ $\mu$s, $T_{\phi}=50$ $\mu$s.]
@@ E.1 Optimized 20-ns pulse for $\hat{X}_{\pi/2}$ gate
E.2 Comparison of pulse robustness with recent literature
Figure 9: Comparison of optimization results between Detuning-robust (DR) and QOC for the $\hat{X}{\pi}$ gate. (a) Average infidelities of a closed system [Eq. (12)] versus pulse durations derived from DR (red) and QOC (blue). The optimal duration selected for DR from the range considered is 20 ns (indicated by a red star). (b) $I$ and $Q$ quadratures of the DR pulse envelopes for the selected 20-ns pulse from (a). (c) Frequency components of the DR pulse envelopes corresponding to the 20-ns pulse. (d) Comparative analysis of the 20-ns pulses from both QOC (blue) and DR (red) in terms of their infidelities as a function of detuning. The average and standard deviation of the infidelities are tabulated considering a uniform distribution of $\rho(\delta)$ for simplicity. [Parameters: $\delta_{i}/2\pi\in{\pm 1,\pm 3,\pm 4,\pm 7,\pm 10}$ MHz, $K\text{q}/2\pi=-200$ MHz, $T_{1}=50$ $\mu$s, $T_{\phi}=50$ $\mu$s.]
Here, we compare the robustness of the pulses derived from our optimization procedure with selected recent literature. First, we obtain robust pulses using the Q-CTRL package [62, 100]. To ensure a fair comparison with our optimization protocol, which aims to generate ancilla pulses robust to frequency changes caused by dispersive shifts from coupled resonators, we adopt the same error Hamiltonian as in Sec. III.A.1. Specifically, the noise operator in Q-CTRL is implemented as a detuning term $\delta\hat{q}^{\dagger}\hat{q}$ with $\delta/2\pi=10$ MHz, matching the maximum value considered in our protocol. Additionally, we apply the same bandwidth constraints as in our optimization, such as a 500 MHz bandwidth for a 20 ns pulse. To differentiate these results from those produced by our optimization procedure, we use the label “Detuning-robust” for the former and “QOC” for the latter. The averaged infidelities [Eq. (12)] of the closed system as a function of pulse duration are shown in Fig. 9(a). While the average infidelities for Detuning-robust (red) and QOC (blue) are similar for shorter durations, QOC demonstrates orders of magnitude lower infidelity for longer durations. To illustrate the robustness details of the Detuning-robust results, we examine the 20-ns case, which exhibits the lowest infidelity. The $I$ and $Q$ quadratures of the pulse envelopes and their corresponding Fourier components are shown in Fig. 9(b) and (c). Note that although the bandwidth constraint is the same for both Detuning-robust and QOC optimizations, the pulse optimized by Detuning-robust exhibits a smaller bandwidth compared to the one obtained from QOC. The infidelity as a function of detuning is presented in Fig. 9(d). For a closed system, the Detuning-robust results (red dashed line) show lower infidelity near resonance but exhibit higher infidelity off-resonance ($\delta/2\pi>5$ MHz) compared to the QOC results (blue dashed line). Notably, in the presence of ancilla decoherence ($T_{1}=50$ $\mu$s, $T_{\phi}=50$ $\mu$s), the infidelity at small detuning is dominated by incoherent error (solid lines). This results in QOC having an overall lower infidelity within the considered range of detuning for robust pulses.
Furthermore, we evaluate our pulse robustness against those reported in Ref. 67. For a fair comparison, we adopt the same parameters: transmon anharmonicity $K_{\text{q}}/2\pi=-260$ MHz, detuning range $\mathbb{D}/2\pi=[-4,2]$ MHz, durations for $\hat{X}_{\pi}$ and $\hat{X}_{\pi/2}$ gates of 70 ns and 80 ns, respectively, and no decoherence error. The comparison is illustrated in Fig. 10, where the blue and red curves represent our results and those from Ref. 67, respectively. Within the chosen range of detuning, our results consistently exhibit significantly lower infidelity.
Figure 10: Infidelity of $\hat{X}_{\pi}$ and $\hat{X}_{\pi/2}$ as a function of detuning. The blue curves plot results obtained from the developed QOC pulses, while the red ones represent infidelities reported in Ref. 67. Note that the data below the gray line are not available in the reference. [Parameters: $K_{\text{q}}/2\pi=-260\,\text{MHz}$, durations for $\hat{X}_{\pi}$ and $\hat{X}_{\pi/2}$ gates of 70 ns and 80 ns, respectively.]
Appendix F Limits of the virtual phase correction in Eq. (22)
Figure 11: Infidelities of $\hat{X}_{\pi}$ gate acting on an ancilla dispersively coupled to a cavity mode. Blue and red curves are realized with Gaussian and DRAG pulses, respectively. The horizontal axis denotes the dispersive shift $\chi$. Solid and dashed lines show the infidelities achieved with and without the cavity phase corrections in Eq. (22), respectively. For reference, infidelities obtained without relative cavity phase are displayed using discrete markers. [Parameters: $K_{\text{q}}/2\pi=-200\,\text{MHz}$, $d_{\text{c}}=10$, with other nonlinearities derived from Eq. (26).]
In Sec. III.2, we demonstrate that the direct application of the robust pulse developed for an isolated ancilla to a composite cavity-ancilla system results in a small reduction in fidelity, which can be almost entirely recovered by applying the virtual phase correction outlined in Eq. (22). It is important to emphasize that this correction is effective only under the condition that the detunings are within the range of robust control, i.e., $\mathcal{I}_{\text{c}}(\delta)\approx 0$ in Eq. (20). To illustrate the limitations of this correction, we present a scenario where the above condition is not satisfied. Specifically, we calculate the infidelity of the $\hat{X}_{\pi}$ gate acting on a cavity-ancilla system, realized with both Gaussian and DRAG pulses. In Fig. 11, the ideal cases with no relative phase associated with the cavity mode are shown as circular dots. In practice, when the relative phase is accounted for, the infidelities display deviations from the ideal cases, as shown by the dashed curves. Unlike the situation with optimized pulses shown in Fig. 5, the infidelities here barely change when the phase corrections are applied (solid curves). This behavior is expected due to the weak robustness of the Gaussian and DRAG pulses, as indicated in Fig. 4. Therefore, the virtual phase correction is only effective when the detuning is within the range of robustness and is primarily compatible with pulses obtained from optimal control.